Solve quadratic equation x2 - x - 1 = 0


1
Expert's answer
2020-08-10T13:59:12-0400

Solution


The problem can be solved using quadratic formula which state: x = "\frac{-b\pm\sqrt{b^2-4ac}}{2a}"

the formula is for equations such as "ax^2+ bx + c = 0"

where in this problem x2−x−1= 0 

a=1, b=-1, c=-1

substituting a,b,and c in the quadratic formula we get;

x ="\frac{-(-1)\pm\ \ \sqrt{(-1)2-4(1)(-1)}}{2\times1}"

this gives out

x = "\frac{1\pm\ \ \sqrt{1\ +\ 4}}{2}"

therefore x ="\frac{1-\ \ \sqrt5}{2} or \frac{1\ +\ \ \sqrt5}{2}"

Hence x=−0.62 or x=1.62


Answer.

x=−0.62 or x=1.62